OpenStudy (anonymous):

Ginger's baby weighed only 3 3/8 pounds when it was born. The doctor said it could not leave the hospital until it weighed 5 1/2 pounds. How many pounds must the baby gain before it can go home to Ginger ?

OpenStudy (anonymous):

5 1/2 - 3 3/8 = 11/2 -27/8 = 44/8 -27/8 = 17/8 pounds